# Construction and classification of differential symmetry breaking operators for principal series representations of the pair $(SO_{0}(4,1),SO_{0}(3,1))$ for special parameters

### v1

@@ Abstract
We construct and give a complete classification of all the differential symmetry breaking operators $\mathbb{D}_{\lambda,\nu}^{N,m}:C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1})\rightarrow C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$, between the spaces of smooth sections of a vector bundle of rank $12N+$ 1 over the $3$-sphere $\mathcal{V}_{\lambda}^{2N+1}\rightarrow S^{3}$, and a line bundle over the $2$-sphere $\mathcal{L}_{m,\nu}\rightarrow S^{2}$ in the special case $|m|=N$.
@@ 3. $\bullet$
In the case $N=m=0$, Problems A and B above are known to be completely solved (see, for instance [7, 13]). When $N=1$, a solution to Problems A and B is given in [12] for $m=0$ (see [22, Rem 1.5]) and in [22] for $|m|\geq 1$. Therefore, the problems above have been completely solved when $N=0,1$ and $m\in{\mathbb{Z}}$.
@@
At the end, we show the existence of a duality between the cases $m$ and $-m$ when $|m|\geq N$ (see Proposition 8.2), proving that the solution to Problems A and B for $m\leq-N$ is equivalent to that for $m>N$; therefore, it suffices to solve one of the cases to obtain the whole solution.
Throughout the text, we use the following notation: :=${\mathbb{N}}:=\{0,1,2,\ldots\}$, +:=${\mathbb{N}}_{+}:=\{1,2,3,\ldots\}$.
@@ 1.1 Organization of the paper
The paper is divided in 9 sections (including one appendix), and is structured as follows. In Section 1.2 we state the main results (Theorems 1.2 and 1.3) and do some remarks. In Section 2 we do a quick review of the F-method in a general setting, the machinery used to prove the main theorems. In Section 3 we specialize the setting to our case $(G,G^{\prime},X,Y)=(SO_{0}(4,1),SO_{0}(3,1),S^{3},S^{2})$, giving detail in the realizations and fixing notations. Sections 4 and 5 are devoted to apply the F-method in two steps, which we call Step 1 and Step 2. By using the main result obtained in Step 2 (Theorem 5.1), we prove in Section 5.1 the main theorems of the paper for $m=N$. After that, we focus on proving Theorem 5.1, which can be divided into two results: Theorems 5.2 and 5.3. These are proved in Sections 6 and 7 respectively. Section 8 shows the existence of a duality between the cases $m$ and $-m$ for $|m|\geq N$, and allows us to prove the main theorems for $m=-N$ by using the obtained results for $m=N$. The appendix at the end collects some properties of Gegenbauer polynomials we use throughout the paper, specially in Section 7.
@@ 1.2 Main results
Suppose $\nu-\lambda\in{\mathbb{N}}$ and let ~$\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ denote the following scalar-valued differential operator (cf. [12, Eq. (2.22)])
~$\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ $=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-4\frac{\partial^{2}}{\partial z\partial\overline{z}},\frac{\partial}{\partial x_{3}}\right),$ (1.2)
@@
Let $\lambda,\nu\in{\mathbb{C}}$ with $\nu-\lambda\in{\mathbb{N}}$, and let $m=\pm N$. Then, any differential symmetry breaking operator in (1.1) is proportional to the differential operator $\mathbb{D}_{\lambda,\nu}^{N,N}$ of order $\nu-\lambda$ given as follows:
@@
Both Theorems 1.2 and 1.3 will be proved in Section 5.1 for $m=N$ and in Section 8 for $m=-N$.
@@
As we pointed out in the introduction, the case $N=|m|=1$ is done in [22]. Concretely, the operators $\mathbb{D}_{\lambda,\nu}^{\pm 1}\equiv\mathbb{D}_{\lambda,\nu}^{1,\pm 1}$ appearing in [22, Eqs. (1.6)–(1.9)] are as follows:
@@ 2 Review of the F-method
$A\rightarrow{\mathbb{C}}^{\times},\enspace a\mapsto a^{\lambda}:=e^{\langle\lambda,\log a\rangle}.$
@@
$\psi\circ\sigma_{\lambda}(\ell)=\tau_{\nu}(\ell)\circ\text{{Ad}}_{\#}(\ell)\psi\quad\forall\enspace\ell\in L^{\prime},$ (2.1)
@@
$\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{W}\right)\psi=0\quad\forall\enspace C\in\mathfrak{n}_{+}^{\prime}.$ (2.2)
@@
[…]
@@
  • Step 2. Solve the equation $\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{W}\right)\psi=0,\quad\forall\enspace C\in\mathfrak{n}_{+}^{\prime}$.
@@ 3 Principal Series Representations of $G=SO_{0}(4,1)$ and $G^{\prime}=SO_{0}(3,1)$
$O(4,1)=\{g\in GL(5,{\mathbb{R}}):Q_{4,1}(gx)=Q_{4,1}(x)\enspace\forall x\in{\mathbb{R}}^{5}\},$
@@
$H_{0}$ $:=E_{1,5}+E_{5,1},$ (3.1)
$C_{j}^{+}$ $:=-E_{1,j+1}+E_{j+1,1}-E_{5,j+1}-E_{j+1,5}$ $(1\leq j\leq 3),$
$C_{j}^{-}$ $:=-E_{1,j+1}+E_{j+1,1}+E_{5,j+1}+E_{j+1,5}$ $(1\leq j\leq 3),$
$X_{p,q}$ $:=E_{q+1,p+1}-E_{p+1,q+1}$ $(1\leq p\leq q\leq 3).$
Here, $E_{p,q}\enspace(1\leq p,q\leq 5)$ represent the elementary $5\times 5$ matrix with a $1$ in the $(p,q)$-entry and zeros elsewhere. One clearly has
@@
$\mathfrak{n}_{+}({\mathbb{R}})$ $:=\ker(\text{{ad}}(H_{0})-\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{n}_{-}({\mathbb{R}})$ $:=\ker(\text{{ad}}(H_{0})+\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{l}({\mathbb{R}})$ $:=\ker\text{{ad}}(H_{0})\subset\mathfrak{g}({\mathbb{R}}).$
@@
and observe that the projection $p:\Xi\rightarrow S^{3},\enspace x\mapsto p(x):=\frac{1}{x_{4}}{}^{t}(x_{0},x_{1},x_{2},x_{3})$ determines a bijection $\widetilde{p}:\Xi/{\mathbb{R}}^{\times}\stackrel{{\scriptstyle\mathclap{\thicksim}}}{{\longrightarrow}}S^{3}$. Let $G:=SO_{0}(4,1)$ be the identity component of $O(4,1)$ and define
@@
$\begin{gathered}M=\left\{\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}:g\in SO(3)\right\}\simeq SO(3),\\[6.0pt] A=\exp({\mathbb{R}}H_{0})\simeq SO_{0}(1,1),\quad N_{+}=\exp(\mathfrak{n}_{+}({\mathbb{R}}))\simeq{\mathbb{R}}^{3}.\end{gathered}$
@@
$\begin{gathered}n_{-}(x):=\exp\left(\sum_{j=1}^{3}x_{j}C_{j}^{-}\right)=I_{3}+\begin{pmatrix}-\frac{1}{2}Q_{3}(x)&-{}^{t}x&-\frac{1}{2}Q_{3}(x)\\ x&0&x\\ \frac{1}{2}Q_{3}(x)&{}^{t}x&\frac{1}{2}Q_{3}(x)\end{pmatrix}.\end{gathered}$ (3.3)
@@
$\mathfrak{n}_{-}({\mathbb{R}})\hookrightarrow G/P\simeq S^{3},\enspace Z\mapsto\exp(Z)\cdot P.$
@@
$(B,\exp(tH_{0}))$ $\longmapsto e^{\lambda t}\sigma^{2N+1}(B),$
@@
[…]
@@
$\psi\circ\sigma_{\lambda}^{2N+1}(\ell)=\tau_{m,\nu}(\ell)\circ\text{{Ad}}_{\#}(\ell)\psi,\quad\forall\enspace\ell\in L^{\prime}=M^{\prime}A,$ (3.9)
$\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{{\mathbb{C}}_{m,\nu}}\right)\psi=0,\quad\forall\enspace C\in\mathfrak{n}_{+}^{\prime}.$ (3.10)
@@
$\enspace(g,\exp(tH_{0}),p(\cdot))$ $\longmapsto p\left(e^{-t}\begin{pmatrix}g&\\ &1\end{pmatrix}^{-1}\cdot\right).$
@@ 4.1 Finite-Dimensional Representation Theory of SO(3)
$\begin{gathered}\theta^{k}:SU(2)\longrightarrow GL_{\mathbb{C}}\left(V^{k+1}\right),\quad g\longmapsto\left(p(\cdot)\mapsto p\left(g^{-1}\cdot\right)\right),\end{gathered}$
@@
$\begin{gathered}\varpi(U(x,y))=\begin{pmatrix}\text{{Re}}(x^{2}-y^{2})&\text{{Im}}(x^{2}+y^{2})&-2\text{{Re}}(xy)\\[3.0pt] -\text{{Im}}(x^{2}-y^{2})&\text{{Re}}(x^{2}+y^{2})&2\text{{Im}}(xy)\\[3.0pt] 2\text{{Re}}(x\overline{y})&2\text{{Im}}(x\overline{y})&|x|^{2}-|y|^{2}\end{pmatrix},\end{gathered}$ (4.3)
@@
$\begin{gathered}\sigma^{2N+1}:SO(3)\longrightarrow GL_{\mathbb{C}}\left(V^{2N+1}\right),\quad B\longmapsto\left(p(\cdot)\mapsto p\left(\left(\widetilde{\varpi}^{-1}(B)\right)^{-1}\cdot\right)\right),\end{gathered}$
@@
denote the space of harmonic polynomials of degree $k$ on ${\mathbb{C}}^{d}$. When $d=2$, this space has dimension at most $2$ and can be explicitly described as follows (see, for instance [6, Thm. 3.1]):
@@
$\begin{array}[]{l l}h_{m+\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m+\ell}({\mathbb{C}}^{2}),&\enspace u_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m+\ell},\\[7.0pt] h_{m-\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m-\ell}({\mathbb{C}}^{2}),&\enspace u_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m-\ell},\\[7.0pt] h_{-m-\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m-\ell}({\mathbb{C}}^{2}),&\enspace u_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m-\ell},\\[7.0pt] h_{-m+\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m+\ell}({\mathbb{C}}^{2}),&\enspace u_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m+\ell}.\end{array}$ (4.7)
@@
[…]
@@
$(V^{2N+1})^{\vee}\otimes{\mathbb{C}}_{m}\otimes\mathcal{H}^{0}({\mathbb{C}}^{2})$ $\simeq\bigoplus_{\ell=1}^{N}{\mathbb{C}}_{m-\ell}\oplus{\mathbb{C}}_{m}\oplus\bigoplus_{\ell=1}^{N}{\mathbb{C}}_{m+\ell}\enspace(k=0),$ (4.10)
$\begin{split}(V^{2N+1})^{\vee}\otimes{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})&\simeq\bigoplus_{\ell=1}^{N}({\mathbb{C}}_{m+k-\ell}\oplus{\mathbb{C}}_{m-k-\ell})\oplus{\mathbb{C}}_{m+k}\oplus{\mathbb{C}}_{m-k}\\ &\quad\oplus\bigoplus_{\ell=1}^{N}({\mathbb{C}}_{m+k+\ell}\oplus{\mathbb{C}}_{m-k+\ell})\enspace(k>0).\end{split}$ (4.11)
The equivalence (i) $\Leftrightarrow$ (iii) follows from taking $SO(2)$-invariants in the decompositions above. In order to find the dimension of $\text{{Hom}}_{SO(2)}\left(V^{2N+1},{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})\right)$ we make use of the condition $|m|\geq N$. If $k=0\in K_{N,m}$, then necessarily $m=\pm N$. In this case there is only one term that does not vanish when taking $SO(2)$-invariants in (4.10), which is ${\mathbb{C}}_{m-N}$ if $m=-N$ and ${\mathbb{C}}_{m+N}$ if $m=N$. Similarly, if $k>0$ we note that the number of terms in (4.11) that are subject to remain when taking $SO(2)$-invariants is $12N+$ 1, one for each value of $k\in K_{N,m}$. But since $|m|\geq N$, we have #$\#K_{N,m}=2N+1$, so only one term remains for each $k\in K_{N,m}$. Therefore $\dim_{\mathbb{C}}\text{{Hom}}_{SO(2)}\left(V^{2N+1},{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})\right)=1$.
@@
The sign index $\pm$ in $h_{k}^{\pm}$ is chosen to coincide with that of m±plus-or-$m\pm\ell$, being $+$ if $m\geq N$ and -- if $m\leq-N$. This choice also matches with the sign of $(\zeta_{1}\pm i\zeta_{2})$.
@@
$A\times\text{{Pol}}^{a}(\mathfrak{n}_{+})\rightarrow\text{{Pol}}^{a}(\mathfrak{n}_{+}),\enspace(\exp(tH_{0}),p)\mapsto e^{-ta}p.$
@@
where $a:=\nu-\lambda$ and the polynomials $g_{k}(t)\enspace(k=0,\dots,2N)$ are given as follows:
@@
[…]
@@
[…]

### v2

@@ Abstract
We construct and give a complete classification of all the differential symmetry breaking operators $\mathbb{D}_{\lambda,\nu}^{N,m}:C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1})\rightarrow C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$, between the spaces of smooth sections of a vector bundle of rank $2N+1$ over the $3$-sphere $\mathcal{V}_{\lambda}^{2N+1}\rightarrow S^{3}$, and a line bundle over the $2$-sphere $\mathcal{L}_{m,\nu}\rightarrow S^{2}$ in the special case $|m|=N$.
@@ 3. $\bullet$
In the case $N=m=0$, Problems A and B above are known to be completely solved (see, for instance [7, 13]). When $N=1$, a solution to Problems A and B is given in [12] for $m=0$ (see [21, Rem 1.5]) and in [21] for $|m|\geq 1$. Therefore, the problems above have been completely solved when $N=0,1$ and $m\in{\mathbb{Z}}$.
@@
At the end, we show the existence of a duality between the cases $m$ and m-m when $|m|\geq N$ (see Proposition 8.2), proving that the solution to Problems A and B for $m\leq-N$ is equivalent to that for $m>N$; therefore, it suffices to solve one of the cases to obtain the whole solution.
Throughout the text, we use the following notation: ${\mathbb{N}}:=\{0,1,2,\ldots\}$, ${\mathbb{N}}_{+}:=\{1,2,3,\ldots\}$.
@@ 1.1 Organization of the paper
The paper is divided in 9 sections (including one appendix), and is structured as follows. In Section 1.2 we state the main results (Theorems 1.2 and 1.3) and do some remarks. In Section 2 we do a quick review of the F-method in a general setting, the machinery used to prove the main theorems. In Section 3 we specialize the setting to our case $(G,G^{\prime},X,Y)=(SO_{0}(4,1),SO_{0}(3,1),S^{3},S^{2})$, giving detail in the realizations and fixing notations. Sections 4 and 5 are devoted to apply the F-method in two steps, which we call Step 1 and Step 2. By using the main result obtained in Step 2 (Theorem 5.1), we prove in Section 5.1 the main theorems of the paper for $m=N$. After that, we focus on proving Theorem 5.1, which can be divided into two results: Theorems 5.2 and 5.3. These are proved in Sections 6 and 7 respectively. Section 8 shows the existence of a duality between the cases $m$ and m-m for $|m|\geq N$, and allows us to prove the main theorems for m=Nm=-N by using the obtained results for $m=N$. The appendix at the end collects some properties of Gegenbauer polynomials we use throughout the paper, specially in Section 7.
@@ 1.2 Main results
Suppose $\nu-\lambda\in{\mathbb{N}}$ and let $\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ denote the following scalar-valued differential operator (cf. [12, Eq. (2.22)])
$\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ $=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-4\frac{\partial^{2}}{\partial z\partial\overline{z}},\frac{\partial}{\partial x_{3}}\right),$ (1.2)
@@
Let $\lambda,\nu\in{\mathbb{C}}$ with $\nu-\lambda\in{\mathbb{N}}$, and let $m=\pm N$. Then, any differential symmetry breaking operator in (1.1) is proportional to the differential operator $\mathbb{D}_{\lambda,\nu}^{N,m}$ of order $\nu-\lambda$ given as follows:
@@
Both Theorems 1.2 and 1.3 will be proved in Section 5.1 for $m=N$ and in Section 8 for m=Nm=-N.
@@
As we pointed out in the introduction, the case $N=|m|=1$ is done in [21]. Concretely, the operators $\mathbb{D}_{\lambda,\nu}^{\pm 1}\equiv\mathbb{D}_{\lambda,\nu}^{1,\pm 1}$ appearing in [21, Eqs. (1.6)–(1.9)] are as follows:
@@ 2 Review of the F-method
$A\rightarrow{\mathbb{C}}^{\times},\kern 4.62497pta\mapsto a^{\lambda}:=e^{\langle\lambda,\log a\rangle}.$
@@
$\psi\circ\sigma_{\lambda}(\ell)=\tau_{\nu}(\ell)\circ\text{{Ad}}_{\#}(\ell)\psi\quad\forall\kern 4.62497pt\ell\in L^{\prime},$ (2.1)
@@
$\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{W}\right)\psi=0\quad\forall\kern 4.62497ptC\in\mathfrak{n}_{+}^{\prime}.$ (2.2)
@@
$\text{{Diff}}_{G^{\prime}}\left(C^{\infty}(X,\mathcal{V}),C^{\infty}(Y,\mathcal{W})\right)$$\text{{Sol}}(\mathfrak{n}_{+};\sigma_{\lambda},\tau_{\nu}).$$\scriptstyle{\displaystyle{{}_{\text{{Symb}}\otimes\text{{id}}}}}\scriptstyle{\thicksim}$
@@
  • Step 2. Solve the equation $\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{W}\right)\psi=0,\quad\forall\kern 4.62497ptC\in\mathfrak{n}_{+}^{\prime}$.
@@ 3 Principal Series Representations of $G=SO_{0}(4,1)$ and $G^{\prime}=SO_{0}(3,1)$
$O(4,1)=\{g\in GL(5,{\mathbb{R}}):Q_{4,1}(gx)=Q_{4,1}(x)\kern 4.62497pt\forall x\in{\mathbb{R}}^{5}\},$
@@
$H_{0}$ $=E_{1,5}+E_{5,1},$ (3.1)
$C_{j}^{+}$ $=-E_{1,j+1}+E_{j+1,1}-E_{5,j+1}-E_{j+1,5}$ $(1\leq j\leq 3),$
$C_{j}^{-}$ $=-E_{1,j+1}+E_{j+1,1}+E_{5,j+1}+E_{j+1,5}$ $(1\leq j\leq 3),$
$X_{p,q}$ $=E_{q+1,p+1}-E_{p+1,q+1}$ $(1\leq p\leq q\leq 3).$
Here, $E_{p,q}\kern 4.62497pt(1\leq p,q\leq 5)$ represent the elementary $5\times 5$ matrix with a $1$ in the $(p,q)$-entry and zeros elsewhere. One clearly has
@@
$\mathfrak{n}_{+}({\mathbb{R}})$ $=\ker(\text{{ad}}(H_{0})-\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{n}_{-}({\mathbb{R}})$ $=\ker(\text{{ad}}(H_{0})+\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{l}({\mathbb{R}})$ $=\ker\text{{ad}}(H_{0})\subset\mathfrak{g}({\mathbb{R}}).$
@@
[…]
@@
$\begin{gathered}M=\left\{\begin{pmatrix}1&&\\ &g&\\ &&1\end{pmatrix}:g\in SO(3)\right\}\simeq SO(3),\\[6.0pt] A=\exp({\mathbb{R}}H_{0})\simeq SO_{0}(1,1),\qquad N_{+}=\exp(\mathfrak{n}_{+}({\mathbb{R}}))\simeq{\mathbb{R}}^{3}.\end{gathered}$
@@
[…]
@@
$\mathfrak{n}_{-}({\mathbb{R}})\hookrightarrow G/P\simeq S^{3},\kern 4.62497ptZ\mapsto\exp(Z)\cdot P.$
@@
$\hskip-5.69046pt(B,\exp(tH_{0}))$ $\longmapsto e^{\lambda t}\sigma^{2N+1}(B),$
@@
$\text{{Diff}}_{G^{\prime}}\left(C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1}),C^{\infty}(S^{2},\mathcal{L}_{m,\nu})\right)$$\text{{Sol}}(\mathfrak{n}_{+};\sigma_{\lambda}^{2N+1},\tau_{m,\nu}),$$\scriptstyle{\displaystyle{{}_{\text{{Symb}}\otimes\text{{id}}}}}\scriptstyle{\thicksim}$ (3.7)
@@
$\psi\circ\sigma_{\lambda}^{2N+1}(\ell)=\tau_{m,\nu}(\ell)\circ\text{{Ad}}_{\#}(\ell)\psi,\quad\forall\kern 4.62497pt\ell\in L^{\prime}=M^{\prime}A,$ (3.9)
$\left(\widehat{d\pi_{\mu}}(C)\otimes\text{{id}}_{{\mathbb{C}}_{m,\nu}}\right)\psi=0,\quad\forall\kern 4.62497ptC\in\mathfrak{n}_{+}^{\prime}.$ (3.10)
@@
$\kern 4.62497pt(g,\exp(tH_{0}),p(\cdot))$ $\longmapsto p\left(e^{-t}\begin{pmatrix}g&\\ &1\end{pmatrix}^{-1}\cdot\right).$
@@ 4.1 Finite-Dimensional Representation Theory of $SO(3)$
$\begin{gathered}\theta^{k}:SU(2)\longrightarrow GL_{\mathbb{C}}\left(V^{k+1}\right),\qquad g\longmapsto\left(p(\cdot)\mapsto p\left(g^{-1}\cdot\right)\right),\end{gathered}$
@@
[…]
@@
$\begin{gathered}\sigma^{2N+1}:SO(3)\longrightarrow GL_{\mathbb{C}}\left(V^{2N+1}\right),\qquad B\longmapsto\left(p(\cdot)\mapsto p\left(\left(\widetilde{\varpi}^{-1}(B)\right)^{-1}\cdot\right)\right),\end{gathered}$
@@
denote the space of harmonic polynomials, homogeneous of degree $k$ on ${\mathbb{C}}^{d}$. When $d=2$, this space has dimension at most $2$ and can be explicitly described as follows (see, for instance [6, Thm. 3.1]):
@@
$\begin{array}[]{l l}h_{m+\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m+\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m+\ell},\\[7.0pt] h_{m-\ell}^{+}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{m-\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}+i\zeta_{2})^{m-\ell},\\[7.0pt] h_{-m-\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m-\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N+\ell}=\xi_{1}^{N-\ell}\xi_{2}^{N+\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m-\ell},\\[7.0pt] h_{-m+\ell}^{-}:V^{2N+1}\longrightarrow{\mathbb{C}}_{m}\otimes\mathcal{H}^{-m+\ell}({\mathbb{C}}^{2}),&\kern 4.62497ptu_{N-\ell}=\xi_{1}^{N+\ell}\xi_{2}^{N-\ell}\longmapsto 1\otimes(\zeta_{1}-i\zeta_{2})^{-m+\ell}.\end{array}$ (4.7)
@@
$\simeq\bigoplus_{\ell=1}^{N}{\mathbb{C}}_{\ell}\oplus{\mathbb{C}}_{0}\oplus\bigoplus_{\ell=1}^{N}{\mathbb{C}}_{-\ell}\qquad\text{ as }SO(2)\text{-modules.}$
@@
$(V^{2N+1})^{\vee}\otimes{\mathbb{C}}_{m}\otimes\mathcal{H}^{0}({\mathbb{C}}^{2})$ $\simeq\bigoplus_{\ell=1}^{N}{\mathbb{C}}_{m-\ell}\oplus{\mathbb{C}}_{m}\oplus\bigoplus_{\ell=1}^{N}{\mathbb{C}}_{m+\ell}\kern 4.62497pt(k=0),$ (4.10)
$\begin{split}(V^{2N+1})^{\vee}\otimes{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})&\simeq\bigoplus_{\ell=1}^{N}({\mathbb{C}}_{m+k-\ell}\oplus{\mathbb{C}}_{m-k-\ell})\oplus{\mathbb{C}}_{m+k}\oplus{\mathbb{C}}_{m-k}\\ &\qquad\oplus\bigoplus_{\ell=1}^{N}({\mathbb{C}}_{m+k+\ell}\oplus{\mathbb{C}}_{m-k+\ell})\kern 4.62497pt(k>0).\end{split}$ (4.11)
The equivalence (i) $\Leftrightarrow$ (iii) follows from taking $SO(2)$-invariants in the decompositions above. In order to find the dimension of $\text{{Hom}}_{SO(2)}\left(V^{2N+1},{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})\right)$ we make use of the condition $|m|\geq N$. If $k=0\in K_{N,m}$, then necessarily $m=\pm N$. In this case there is only one term that does not vanish when taking $SO(2)$-invariants in (4.10), which is ${\mathbb{C}}_{m-N}$ if m=Nm=-N and ${\mathbb{C}}_{m+N}$ if $m=N$. Similarly, if $k>0$ we note that the number of terms in (4.11) that are subject to remain when taking $SO(2)$-invariants is $2N+1$, one for each value of $k\in K_{N,m}$. But since $|m|\geq N$, we have $K_{N,m}=2N+1$, so only one term remains for each $k\in K_{N,m}$. Therefore $\dim_{\mathbb{C}}\text{{Hom}}_{SO(2)}\left(V^{2N+1},{\mathbb{C}}_{m}\otimes\mathcal{H}^{k}({\mathbb{C}}^{2})\right)=1$.
@@
The sign index $\pm$ in $h_{k}^{\pm}$ is chosen to coincide with that of $m\pm\ell$, being ++ if $m\geq N$ and - if $m\leq-N$. This choice also matches with the sign of $(\zeta_{1}\pm i\zeta_{2})$.
@@
$A\times\text{{Pol}}^{a}(\mathfrak{n}_{+})\rightarrow\text{{Pol}}^{a}(\mathfrak{n}_{+}),\kern 4.62497pt(\exp(tH_{0}),p)\mapsto e^{-ta}p.$
@@
where $a:=\nu-\lambda$ and the polynomials $g_{k}(t)\kern 4.62497pt(k=0,\dots,2N)$ are given as follows:
@@
$\text{{Diff}}_{G^{\prime}}\left(C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1}),C^{\infty}(S^{2},\mathcal{L}_{N,\nu})\right)$$\text{{Sol}}(\mathfrak{n}_{+};\sigma_{\lambda}^{2N+1},\tau_{N,\nu}).$$\scriptstyle{\displaystyle{{}_{\text{{Symb}}\otimes\text{{id}}}}}\scriptstyle{\thicksim}$
@@
By (3.7) and Theorem 5.1, it is enough to compute the inverse of the symbol map $\text{{Symb}}\otimes\text{{id}}$ of $\psi=\sum_{k=0}^{2N}\left(T_{a-k}g_{k}\right)(\zeta)h_{k}^{+}$, where $g_{k}$ are the polynomials given in (5.1) and $h_{k}^{+}$ are defined in (4.7).